OpenStudy (lexy148):

determine the roots and y-intercept of the following function using algebra. g(x)=x^4−2x^3−7x^2+20x−12

OpenStudy (danjs):

what you have so far?

OpenStudy (danjs):

It looks like that will factor nicely, the roots are where the graph crosses the x-axis, or y=g(x)=0

OpenStudy (danjs):

Which way did you learn to factor , the synthetic division ?

OpenStudy (danjs):

have to go for a while sorry, g(x) = (x-2)(x-2)(x-1)(x+3) = 0 those are the factors, gl